# TrackPGD: A White-box Attack using Binary Masks against Robust Transformer Trackers

### v1

@@ Abstract
Object trackers with transformer backbones have achieved robust performance on visual object tracking datasets. However, the adversarial robustness of these trackers has not been well studied in the literature. Due to the backbone differences, the adversarial white-box attacks proposed for object tracking are not transferable to all types of trackers. For instance, transformer trackers such as MixFormerM still function well after black-box attacks, especially in predicting the object binary masks. We are proposing a novel white-box attack named TrackPGD, which relies on the predicted object binary mask to attack the robust transformer trackers. That new attack focuses on annotation masks by adapting the well-known SegPGD segmentation attack, allowing to successfully conduct the white-box attack on trackers relying on transformer backbones. The experimental results indicate that the TrackPGD is able to effectively attack transformer-based trackers such as MixFormerM, OSTrackSTS, and TransT-SEG on several tracking datasets.
@@ 1 Introduction
Adversarial attacks [1, 2, 3, 4] can significantly harm neural networks performances by adding carefully crafted imperceptible noise to the input. For instance, classification networks [5, 6] are highly susceptible to such adversarial perturbations. For object segmentation and object tracking, some recent attacks [7, 1, 2, 4, 3] have been developed to disrupt the object masks and bounding boxes output, respectively. Since 2020, the Visual Object Tracking (VOT) challenge [8] introduced a new sub-challenge that requires object trackers to present binary masks as output and be evaluated based on their mask generation ability. In addition, novel trackers with transformer backbones [9, 10, 11] have emerged, whose approaches infer the object mask, bounding box, and confidence score as a triple output. As a result, object binary masks have become increasingly important both as a valid output of trackers and a valid criterion for evaluation in tracking benchmarks.
Visual tracking can be compromised by adversarial attacks in different ways, by manipulating elements such as object bounding boxes [2], classification labels [3, 4, 1], and regression labels [3, 1]. Like other adversarial attacks, the object tracker’s attacks can be categorized into two categories: white-box attacks [7, 3], where the attacker has access to the model’s gradients, and black-box attacks [2], where only the tracker’s output is available. In Table 1, several attack methods and their attack proxy are listed along with their applicability to some new transformer tracker, MixFormerM [9], OSTrackSTS [10] and TransT-SEG [11].
Method Attack Proxy MixFormerM [9] OSTrackSTS [10] TransT-SEG [11]
BB IoU [2] Object bbox $\checkmark$ $\checkmark$ $\checkmark$
CSA [4] Object bbox, heat-maps $\checkmark$ $\checkmark$ $\checkmark$
WB SPARK [1] Regr. and classif. labels $\times$ $\times$ $\checkmark$
RTAA [3] Regr. and classif. labels $\times$ $\times$ $\checkmark$
TrackPGD Object binary mask $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Characterization of attacks methods on transformer-based trackers. TrackPGD is the only applicable White-Box (WB) attack against MixFormerM, and OSTrackSTS, using a new proxy (i.e. binary mask), while the two other methods that can be applied to these trackers, IoU and CSA, are Black-Box (BB) methods manipulating object bounding box.
@@
Figure 1: The binary masks generated by MixFormerM [9] and OSTrackSTS [10] attacked by the IoU [2] vs. TrackPGD. The green/red masks are the before and after the attack outputs.
Table 1 presents several attacks along with their respective attack proxies and specifies their applicability to new transformer trackers. For instance, MixFormerM [9] does not output the labels of object candidates. Therefore, the attack algorithms with the objects labels proxy, such as RTAA [3] and SPARK [1], are not applicable to MixFormerM as a white-box attack. In this paper, we present a white-box attack, named TrackPGD, to attack the transformer trackers via a new proxy, that is the object binary mask. Figure 1 indicates TrackPGD performance on crashing binary masks generated by MixFormerM [9] and OSTrackSTS [10]. The green/red masks represents the predicted mask before and after the attack, respectively. As a white-box attack applicable on both MixFormerM and OSTrackSTS, the binary mask after the attack (red mask) caused more damage than the other applicable attack, i.e. Intersection over Union (IoU) method [2]. Although the IoU method may degrade the MixFormerM’s bounding box, the binary mask remained almost unchanged. Yet, the bounding box can be obtained from the object binary mask. Furthermore, a new tracking dataset such as VOT2022STS [12] evaluates the tracker’s performance based on only the generated mask, making the target object binary masks key elements to consider in order to decrease the robustness of the tracker. Our contributions can be summarized as follows:
@@
  • MixFormerM experienced significant accuracy ($-75\%$), average overlap ($-97\%$) and robustness ($-91\%$) reductions after applying TrackPGD, evaluated on the VOT2022STS dataset.
@@ 2 Related works
Robust Object Trackers Tracker robustness was initially defined as the frequency of failures, marked by the overlap between prediction and ground truth dropping to zero. This definition is employed in datasets like VOT2016 [13] and VOT2018 [14]. In VOT evaluation for short-term tracking, the tracker is reset after each failure. A newer protocol introduced in VOT2020 [8] includes multiple anchors in each test sequence spaced approximately 50 frames apart to facilitate anchor-based short-term tracking [8, 12]. This protocol measures robustness based on the duration of video subsequences where the overlap falls below a fixed threshold for several frames. In contrast, other tracking datasets, like LaSOT benchmark [15], conduct a one-pass evaluation. The three most robust trackers on the VOT2022STS [12] were the Multi-Scale approach based on Object associating with Transformers (MS-AOT) [16, 17], Distractor-Aware Multi-head Tracker (DAMTMask) [18, 19], and MixFormerM [9]. MS-AOT used MixFormer [9] for targets smaller than $1/900$ frame size and for larger targets, a multi-scale object association approach is employed to perform tracking task. DAMTMask integrates MixFormer [9] and TransT-SEG [11] to construct a robust performance. The MixFormer [9] tracker proposed a Mixed Attention Module (MAM) to jointly extract features of search and template regions. MixFormerM [9, 12] contains two networks, one tracker network and the other is AlphaRefine [20] for object segmentation. But TransT-SEG [11] jointly predicts the object bounding box and generates the object binary mask via one single network that uses a light cross-attention mechanism to fuse deep features. Another transformer-based tracker, OSTrackSTS [10], gained the third rank on the VOT-STS2022 [12] sequestered dataset after MS-AOT and DAMTMask. The OSTrackSTS [10] is the One-Stream Track method [10] coupled with the AlphaRefine [20] to provide object binary masks.
@@
Adversarial attacks for object trackers The attack proxy can be classification labels [3, 4, 1], regression output [3] or even the motion features of the object [2, 3]. In the Robust Tracking against Adversarial Attacks (RTAA) method [3], specific labels for classification and regression are modified. An online incremental attack called SPARK [1] has been developed for object trackers that use the previous frame perturbations and manipulate the object classification and regression labels to generate the adversarial frame. The Cooling-Shrinking Attack (CSA) method, as described in [4], involves training two Generative Adversarial Networks (GANs) to create adversarial images for the SiamRPN++ tracker [21]. The cooling term cools down the heated pixels in the heatmap of SiamRPN++, while the shrinking term shrinks the predicted bounding box. The IoU attack is a black-box approach that uses the predicted bounding box to inject two types of noises [2] into frames. The directional and orthogonal perturbations are injected in a given frame to degrade the predicted bounding box from the previous prediction.
3 Method
3.1 Background: SegPGD
@@
SegPGD [7] loss for a segmentation network ($L_{\text{SegPGD}}$) is defined to steer the neural network in providing an inaccurate segmentation map. This loss is computed between the mask annotation $Y$ and the adversarial segmentation map $f_{\text{seg}}(X^{\text{adv}})$ generated by the segmentation network from the adversarial image $X^{\text{adv}}$:
@@
where $\lambda$ is a coefficient that dynamically changes over iterations and $H$ and $W$ are the frame’s height and width, respectively. Here, $L_{\text{pos}}$ and $L_{\text{neg}}$ denote the cross-entropy loss restricted to correctly classified and wrongly classified pixels, respectively. The motivation for this reweighting stems from the observation that gradients of cross-entropy for wrongly classified examples tend to have larger amplitude, while the attack aims primarily at correctly classified pixels [7].
To adapt SegPGD for use against object trackers and to target the predicted binary masks, we must overcome two key challenges. Firstly, while SegPGD operates on multi-class segmentation networks, binary masks only have two classes, necessitating the switching of classification labels between pixels. Secondly, the labels in the majority of samples are imbalanced, with the number of object pixels being significantly fewer than background pixels. To address these hurdles and effectively attack trackers, we introduce TrackPGD.
3.2 Proposed method: TrackPGD
Let us formalize object tracking as using a tracker $\mathcal{F}(\cdot)$ that processes a video sequence $\mathcal{V}=\{I_{0},\ldots,I_{L}\}$ with $L+1$ frames, having the first frame object mask $M_{0}$ provided to initialize the tracker. The transformer tracker is expected to estimate the object mask $M_{\tau}$ at each time step $\tau\in\{1,\ldots,L\}$ from the given video frame $I_{\tau}$. The goal of the attack method is thus to create an adversarial video frame $I^{\text{adv}}_{\tau}$ at each time step to mislead the object tracker in its mask prediction. The transformer tracker predicts at each time $\tau$ a mask $M_{\tau}$ which assigns to each pixel a probability of being an object. The TrackPGD attack aims to mislead the tracker in making these assignments by creating the adversarial frame $I^{\text{adv}}_{\tau}$.
For object tracking, the mask annotation $G_{\tau}$ consists of two labels (object (1) and background (0)) and it follows that finding adversarial examples can be achieved at optimization step $t$ both by maximizing $L_{\text{SegPGD}}(M^{t},G_{\tau})$ or by minimizing $L_{\text{SegPGD}}(M^{t},1-G_{\tau})$, where we consider the mask $M_{\tau-1}$ predicted on the previous frame as the ground truth $G_{\tau}$ at step $\tau$. In the TrackPGD framework, we therefore decide to combine these two complementary objectives into a difference loss $L_{\Delta}$ as follows:
@@ 3.2 Proposed method: TrackPGD
Since the object pixels and background pixels are imbalanced in the video frame, we employed the focal loss [22]. It is a useful loss for training highly imbalanced data, which is the case, here, for the object and background pixels. For TrackPGD, instead of using cross-entropy as in the original equation of focal loss [22], we use the $L_{\Delta}$ loss as follows:
@@
where $\mathcal{P}$ is a probability map derived from the application of the sigmoid function on the predicted binary mask $M_{t}$(before thresholding) and $\gamma$ is the focusing parameter. This $\alpha$-balanced variant of the focal loss used is similar to the TransT-SEG [11] except that $L_{\Delta}$ is used instead of cross-entropy. Then, we compute the dice loss [23] between $M^{t}$ and $G_{\tau}$ as follows:
@@
where IoU is the Intersection over Union loss between the ground truth $G_{\tau}$ and the predicted binary mask $M^{t}$. The combination of focal loss (Section 3.2), and dice loss (Eq. 4), is used to compute the TrackPGD loss as follows:
@@
where the $\lambda_{1}$ and $\lambda_{2}$ balance the focal and dice terms. The pseudocode of TrackPGD is presented in Algorithm 1. The sign of step size $\alpha$, line 9 in Algorithm 1, is determined according the best performance of each tracker, following the decoupling of the norm and direction of gradients in adversarial attacks [24].
2:${I}_{\text{adv}}^{0}\leftarrow I_{\tau}$triangleright initialization
4:for $t=1\ldots T$ do
6:     $L_{\Delta}\leftarrow L_{\text{SegPGD}}(M^{t},G_{\tau})-L_{\text{SegPGD}}(M^{t},1-G_{\tau})$
7:     $L_{\text{focal}}\leftarrow\alpha_{t}(1-p_{t})^{\gamma}L_{\Delta}$
8:     $L_{\text{dice}}\leftarrow 1-2\,\text{IoU}(M^{t},G_{\tau})$
10:     ${I}_{\text{adv}}^{t}\leftarrow I_{\text{adv}}^{t-1}+\alpha\,\mathrm{sign}(\nabla_{I_{\text{adv}}^{t-1}}L)$
12:end for

### v2

# TrackPGD: Efficient Adversarial Attack using Object Binary Masks against Robust Transformer Trackers

@@ Abstract
Adversarial perturbations can deceive neural networks by adding small, imperceptible noise to the input. Recent object trackers with transformer backbones have shown strong performance on tracking datasets, but their adversarial robustness has not been thoroughly evaluated. While transformer trackers are resilient to black-box attacks, existing white-box adversarial attacks are not universally applicable against these new transformer trackers due to differences in backbone architecture. In this work, we introduce TrackPGD, a novel white-box attack that utilizes predicted object binary masks to target robust transformer trackers. Built upon the powerful segmentation attack SegPGD, our proposed TrackPGD effectively influences the decisions of transformer-based trackers. Our method addresses two primary challenges in adapting a segmentation attack for trackers: limited class numbers and extreme pixel class imbalance. TrackPGD uses the same number of iterations as other attack methods for tracker networks and produces competitive adversarial examples that mislead transformer and non-transformer trackers such as MixFormerM, OSTrackSTS, TransT-SEG, and RTS on datasets including VOT2022STS, DAVIS2016, UAV123, and GOT-10k.
@@ 1 Introduction
Adversarial attacks (Guo et al., 2020; Jia et al., 2021, 2020; Yan et al., 2020) can significantly harm neural networks performances by adding carefully crafted imperceptible noise to the input. For instance, classification networks are highly susceptible to such adversarial perturbations (Shao et al., 2022; Mahmood et al., 2021). For object segmentation and object tracking, some recent attacks (Gu et al., 2022; Guo et al., 2020; Jia et al., 2021) have been developed to disrupt the object masks and bounding boxes output, respectively. Since 2020, the Visual Object Tracking (VOT) challenge (Kristan et al., 2020) introduced a new sub-challenge that requires object trackers to present binary masks as output and be evaluated based on their mask generation ability. In addition, novel trackers with transformer backbones have emerged (Cui et al., 2022; Ye et al., 2022; Chen et al., 2023; Bhat et al., 2020), whose approaches infer the object mask, bounding box, and confidence score as a triple output. Given that objects with complex shapes are rarely aligned with the frame axis, the corresponding axis-aligned bounding boxes contain more background pixels than object pixels (Paul et al., 2022). Therefore, object bounding boxes are not the best option to describe an object’s appearance in comparison to object masks. As a result, object binary masks have become increasingly important both as a valid output of trackers and a valid criterion for evaluation in tracking benchmarks. Visual tracking can be compromised by adversarial attacks in different ways, by manipulating elements such as object bounding boxes (Jia et al., 2021), classification labels (Jia et al., 2020; Yan et al., 2020; Guo et al., 2020), and regression labels (Jia et al., 2020; Guo et al., 2020). In order to apply white-box attacks on new transformer trackers (Cui et al., 2022; Ye et al., 2022; Chen et al., 2023; Bhat et al., 2020), the adversarial loss should be computed using the tracker’s elements and outputs. While some white-box attacks, such as SPARK (Guo et al., 2020) and RTAA (Jia et al., 2020), are not applicable to new robust object trackers, such as RTS (Paul et al., 2022) and MixFormer (Cui et al., 2022), the black-box attacks are not effective enough to challenge these trackers. In this work, we propose TrackPGD, a novel and efficient white-box attack that uses the object binary mask as a proxy. Figure 1 illustrates TrackPGD’s effectiveness in disrupting binary masks of MixFormerM (Cui et al., 2022) and OSTrackSTS (Ye et al., 2022), with green/red masks representing predictions before/after the attack. TrackPGD caused more damage than the Intersection over Union (IoU) method (Jia et al., 2021), which only slightly degraded the bounding box while leaving the binary mask mostly unchanged. The IoU method is the most recent applicable attack against MixFormerM (Cui et al., 2022), OSTrackSTS (Ye et al., 2022) and RTS (Paul et al., 2022).
@@
Figure 1: The binary masks generated by MixFormerM (Cui et al., 2022) and OSTrackSTS (Ye et al., 2022) attacked by our TrackPGD vs. IoU (Jia et al., 2021) which is the most recent and applicable attack for these two trackers up to now. The green/red masks are the before/after the attack outputs.
Table 1: Characterization of attack methods– IoU (Jia et al., 2021), CSA (Yan et al., 2020), SPARK (Guo et al., 2020), RTAA (Jia et al., 2020) and TrackPGD (ours)– on robust object trackers including MixFormerM (Cui et al., 2022), OSTrackSTS (Ye et al., 2022), TransT-SEG (Chen et al., 2023) and RTS (Paul et al., 2022). TrackPGD is the only applicable White-box attack against MixFormerM, OSTrackSTS and RTS, using a new proxy (i.e. binary mask), while the two other methods that may be applied to these trackers, IoU and CSA, are Black-box methods manipulating bounding box.
Attack Setting Method Attack Proxy MixFormerM OSTrackSTS TransT-SEG RTS
Black-box IoU Object bbox
CSA Object bbox, heat-maps $\times$
White-box SPARK Regression and classification labels $\times$ $\times$ $\times$
RTAA Regression and classification labels $\times$ $\times$ $\times$
TrackPGD Object binary mask
Targeting object binary masks in adversarial attacks is essential due to their increasing use in tracking backbones (Cui et al., 2022; Ye et al., 2022; Yan et al., 2021; Paul et al., 2022) and datasets (Kristan et al., 2023). In Table 1, several attack methods and their attack proxy are listed along with their applicability to some new transformer trackers, MixFormerM (Cui et al., 2022), OSTrackSTS (Ye et al., 2022) and TransT-SEG (Chen et al., 2023). Using a novel attack proxy, the object binary mask, TrackPGD is developed to generate the adversarial frame patches from the predicted binary mask. In addition, the experimental results on object bounding box evaluation demonstrate that the TrackPGD is an effective attack over neural network-based object trackers. Table 1 lists various attacks, the proxies, and applicability to new trackers. For example, MixFormerM (Cui et al., 2022) does not output object candidate labels, making label-based attacks such as RTAA (Jia et al., 2020) and SPARK (Guo et al., 2020) unsuitable as white-box attacks. In a nutshell, our contributions can be summarized as follows:
@@
  • We demonstrate that MixFormerM suffers significant accuracy ($-75\%$), average overlap ($-97\%$) and robustness ($-91\%$) reductions after applying TrackPGD, evaluated on the VOT2022STS dataset.
@@ 2 Related works
Robust Object Trackers Tracker robustness was initially defined as the frequency of failures, marked by the overlap between prediction and ground truth dropping to zero. This definition is employed in datasets like VOT2016 (Kristan et al., 2016) and VOT2018 (Kristan et al., 2018). A newer protocol introduced in VOT2020 (Kristan et al., 2020) includes multiple anchors in each test sequence spaced approximately 50 frames apart to facilitate anchor-based short-term tracking (Kristan et al., 2020, 2023). In contrast, other tracking datasets, like LaSOT benchmark (Fan et al., 2019), conduct a one-pass evaluation. The MixFormerM (Cui et al., 2022) was one of the three most robust trackers on the VOT2022STS (Kristan et al., 2023). The MixFormer (Cui et al., 2022) tracker proposed a Mixed Attention Module (MAM) to jointly extract features of search and template regions. MixFormerM (Cui et al., 2022; Kristan et al., 2023) contains two networks, one tracker network and the other is AlphaRefine (Yan et al., 2021) for object segmentation. But TransT-SEG (Chen et al., 2023) jointly predicts the object bounding box and generates the object binary mask via a single network that uses a light cross-attention mechanism to fuse deep features. Another transformer tracker, OSTrackSTS (Ye et al., 2022), gained the third rank on the VOT-STS2022 (Kristan et al., 2023) sequestered dataset. The OSTrackSTS (Ye et al., 2022) is the One-Stream Track method (Ye et al., 2022) coupled with the AlphaRefine (Yan et al., 2021) to provide object binary masks. The Robust Tracking by Segmentation (Paul et al., 2022) is a new robust discrimintative tracker to infer the object binary mask from a whole video frame. Using a segmentation branch, and an instance localization branch, RTS achieved a robust tracking performance by predicting the target mask in the tracking process (Kristan et al., 2023).
@@
Adversarial attacks for object trackers The attack proxy can be classification labels (Jia et al., 2020; Yan et al., 2020; Guo et al., 2020), regression output (Jia et al., 2020) or even the motion features of the object (Jia et al., 2021, 2020). In the Robust Tracking against Adversarial Attacks (RTAA) method (Jia et al., 2020), specific labels for classification and regression are modified. An online incremental attack called SPARK (Guo et al., 2020) has been developed for object trackers that use the previous frame perturbations and manipulate the object classification and regression labels to generate the adversarial frame. The Cooling-Shrinking Attack (CSA) method, as described in (Yan et al., 2020), involves training two Generative Adversarial Networks (GANs) to create adversarial images for the SiamRPN++ tracker (Li et al., 2019). The cooling term cools down the heated pixels in the heatmap of SiamRPN++, while the shrinking term shrinks the predicted bounding box. The IoU attack is a black-box approach that uses the predicted bounding box to inject two types of noises (Jia et al., 2021) into frames. The directional and orthogonal perturbations are injected in a given frame to degrade the predicted bounding box from the previous prediction.
3 Preliminaries: SegPGD
@@
The SegPGD (Gu et al., 2022) loss for a segmentation network ($L_{\text{SegPGD}}$) is defined to steer the neural network toward providing an inaccurate segmentation map. This loss is computed between the mask annotation $Y$ and the adversarial segmentation map $f_{\text{seg}}(X^{\text{adv}})$ generated by the segmentation network from the adversarial image $X^{\text{adv}}$:
@@
where $\lambda$ is a coefficient that dynamically changes over iterations and $H$ and $W$ are the frame height and width, respectively. Here, $L_{\text{pos}}$ and $L_{\text{neg}}$ denote the cross-entropy loss restricted to correctly classified and wrongly classified pixels, respectively. The motivation for this reweighting stems from the observation that gradients of cross-entropy for wrongly classified examples tend to have larger amplitude, while the attack aims primarily at correctly classified pixels (Gu et al., 2022).
To adapt SegPGD for use against object trackers and to target the predicted binary masks, we must overcome two key challenges. Firstly, while SegPGD operates on multi-class segmentation networks, binary masks only have two classes, necessitating the switching of classification labels between pixels. Secondly, the labels in the majority of samples are imbalanced, with the number of object pixels being significantly fewer than background pixels. To address these hurdles, we introduce TrackPGD as a white-box attack for object trackers which predict binary masks.
4 Proposed method: TrackPGD
Let us formalize object tracking as using a tracker $\mathcal{F}(\cdot)$ that processes a video sequence $\mathcal{V}=\{I_{0},\ldots,I_{L}\}$ with $L+1$ frames, having the first frame object mask $M_{0}$ provided to initialize the tracker. The transformer tracker is expected to estimate the object mask $M_{\tau}$ at each time step $\tau\in\{1,\ldots,L\}$ from the given video frame $I_{\tau}$. The goal of the attack method is thus to create an adversarial video frame $I^{\text{adv}}_{\tau}$ at each time step to mislead the object tracker in its mask prediction. The transformer tracker predicts at each time $\tau$ a mask $M_{\tau}$ which assigns to each pixel a probability of being an object. The TrackPGD attack aims to mislead the tracker in making these assignments by creating the adversarial frame $I^{\text{adv}}_{\tau}$. For object tracking, the mask annotation $G_{\tau}$ consists of two labels (object (1) and background (0)). It follows that finding adversarial examples can be achieved at optimization step $t$ both by maximizing $L_{\text{SegPGD}}(M^{t},G_{\tau})$ or by minimizing $L_{\text{SegPGD}}(M^{t},1-G_{\tau})$, where we consider the mask $M_{\tau-1}$ predicted on the previous frame as the ground truth $G_{\tau}$ at step $\tau$. In the TrackPGD framework, we therefore decide to combine these two complementary objectives into a difference loss $L_{\Delta}$ as follows:
@@ 4 Proposed method: TrackPGD
Since the object pixels and background pixels are imbalanced in the video frame, we employed the focal loss (Lin et al., 2020). It is a useful loss for training highly imbalanced data, which is the case, here, for the object and background pixels. For TrackPGD, instead of using cross-entropy as in the original equation of focal loss (Lin et al., 2020), we use the $L_{\Delta}$ loss as follows:
@@
where $\mathcal{P}$ is a probability map derived from the application of the sigmoid function on the predicted binary mask $M_{t}$ (before thresholding) and $\gamma$ is the focusing parameter. This $\alpha$-balanced variant of the focal loss used is similar to the TransT-SEG (Chen et al., 2023) except that $L_{\Delta}$ is used instead of cross-entropy. Then, we compute the dice loss (Milletari et al., 2016) between $M^{t}$ and $G_{\tau}$ as follows:
@@
where IoU is the Intersection over Union loss between the ground truth $G_{\tau}$ and the predicted binary mask $M^{t}$. The combination of focal loss (Equation 4), and dice loss (Equation 4), is used to compute the TrackPGD loss as follows:
@@
where the $\lambda_{1}$ and $\lambda_{2}$ balance the focal and dice terms.
For a detailed explanation of the TrackPGD algorithm and additional studies, please refer to the Supplementary materials.